A railroad flatcar is travelling to the right at speed of 13 m/s relative to an observer on the ground. Someone is riding a motor scooter on the flatcar. What is the velocity of motor scooter relative to the flatcar if its relative velocity to the observer on the ground is (a) 18 m/s to the right? (b) 3 m/s to the left? (c) zero?


1
Expert's answer
2021-07-12T12:11:47-0400

The velocity of the scooter relative to the flatcar is given as follows (see https://en.wikipedia.org/wiki/Velocity-addition_formula#Galilean_relativity):


"u' = u-v"

where "u" is the velocity of the scooter relative to the observer, and "v = 13m\/s" is the velocity of the flatcar relative to the observer.

Let's choose the direction to the right as a positive direction. Then:


a) "u = 18m\/s".


"u' = 18m\/s-13m\/s = 5m\/s"

b) "u = -3m\/s" (minus because it is directed to the left).


"u' = -3m\/s - 13m\/s = -16m\/s"

c) "u = 0m\/s".


"u' = 0 - 13m\/s = -13m\/s"

Answer. a) 5m/s to the right, b) 16m/s to the left, c) 13m/s to the left.